From GlueXWiki
Jump to: navigation, search

August 18, 2011 FDC meeting


  1. Production Construction Tracking (Dave)
  2. Engineering (Bill)
  3. Electronics update (Chris)
  4. Chamber testing at EEL126 (cell#2) (Beni)
    • Status with the HV
    • Deadened area tests: [15]
  5. Other


Participants: Bill, Dave, Beni, Simon, and Lubomir. We covered important problems and at the same time we were missing people that could help with the discussions and decisions.


- Dave: Working now on the second cathode for cell#6. Another two cathodes are needed (type 1 and 2) to finish the cathodes for the first production package. There was a problem with the cutter; the springs will be replaced. The wire plane #5 is ready, yesterday we did the deadening. Right after the deadening a wire sample was made using the same chemical. The stringing stopped because of problems with the HV caps on the next wire frame #6.

- Beginning with the wire frame #4, we started testing the HV caps before stringing. It requires some conditioning but few days were enough to test wire frame #5. Now we are having problems with wire frame #6. We identified so far 8 caps that need to be replaced. After the meeting we continued with the tests and we found another 4, so 12 in total. Anatoly, who removed the caps, mentioned that on 5 of these there were solder balls below the caps somewhere between the pads (see pictures linked above). The balls are actually flat (like coins). Therefore we suspect that they were there before the caps were un-soldered by Anatoly. If this is the case then we need to consider new procedures for cleaning/testing/replacing the caps on the boards before the lamination.


- Bill showed pictures (linked above) of the deadened area of one of the samples made by the W&M microscope. The surfaces, including the transition areas, look well polished and free of any sharp edges. The thicknesses along the wire and also for the different wires are very similar of about 66-68 microns, except at the very end of the deadened area where it is ~57 microns. Bill will look also at the last wire sample.

- Do we need to cut the gusset ring? As estimated by Eugene in case of magnet trip without the cut we may have up to ~7mV e.m.f. That means in the ring (~10^-3 Ohms) we will have 7A current and the power will by ~50 mWatt. Nobody from the participants of the meeting was aware of any problem this power can cause. At the same time cutting the rings introduces some issues: flatness problems, dust from the fiber glass/carbon.


- Chris (not at the meeting) is testing the daughter boards and out ~200 he found only several problems mostly on the connectors, that were easily fixed.

Chamber testing at 126 (cell #2)

- On the HV sector #4 we found that there is ~2 MOhm between the field and sense wires. There's no connection from both field and sense wires to ground which indicates most likely there's no broken wire. To further investigate the problem we will apply the same potential to both field and sense wires. This was done already and the HV was high enough to see signals with an iron source.

- Beni was doing tests with HV sectors 1,2,3 using external tracking. The chamber trips once per one or two days. Beni studied the deadened area by looking at the chamber efficiency. The plot (page 572) shows the track position on the FDC as reconstructed from the IU chambers if there was a signal from the FDC wires: 24 channels on one side in the middle of the chamber. One can clearly see the deadened area also as a projection on one wire. The drift time spectrum is still a concern (bottom picture) therefore Beni decided to switch to a premixed bottle 89.7/10.3 Ar/CO2.